OpenStudy (anonymous):

find the inverse function of y=3/2(x-3), the answer is y=2x+9/3 steps? also y=+/-2 sqrt(x-3) +3 , the answer is y=x^2 -6x+12 . Please show me the steps how

OpenStudy (ranga):

To find the inverse, 1. Replace x with y and y with x. 2. Solve for y.

OpenStudy (anonymous):

i know that , can u please show me step by step?

OpenStudy (ranga):

y=3/(2x-3) Replace x with y and y with x. What do you get?

OpenStudy (anonymous):

im having difficulty isolating y

OpenStudy (ranga):

Do one step at a time. Replace x with y and y with x. What do you get?

OpenStudy (anonymous):

i've got y= 2x/3 +3,which i snot the answer, i need to know what im doing wrong. can you please just show me your steps to get that answer please

OpenStudy (ranga):

Well you won't even answer any of my basic question. I will let someone else help you.

OpenStudy (anonymous):

right now i have , 2x/3 = 1/y-3 , how can i isolate y?

OpenStudy (jdoe0001):

cross-multiply

OpenStudy (jdoe0001):

\(\bf {\color{red}{ y}}=\cfrac{3}{2{\color{blue}{ x}}-3}\qquad inverse\implies {\color{blue}{ x}}=\cfrac{3}{2{\color{red}{ y}}-3} \\ \quad \\ 2{\color{red}{ y}}-3=\cfrac{3}{{\color{blue}{ x}}}\) add 3 and divide by 2, on both sides

OpenStudy (anonymous):

i;ve made a mistake in the question it is actually y=3/2(x-3)